OpenStudy (anonymous):

A=i+j The value of B=i-j (A+B).(A-B) It is simple but i don't know where i am making a mistake

OpenStudy (vishweshshrimali5):

(A+B)(A-B) = A.A - A.B + B.A - B.B = A.A - B.B = |A|^2 - |B|^2

OpenStudy (vishweshshrimali5):

This is one method in which you will only have to calculate the magnitude of each vector and you would get the answer :)

OpenStudy (vishweshshrimali5):

Another method is to calculate (A+B) and (A-B) and then evaluate the dot product

OpenStudy (vishweshshrimali5):

Now which one did you use ?

OpenStudy (anonymous):

I did A+B =2i A-B=2j so 2i*2j = 4 ?

OpenStudy (vishweshshrimali5):

Actually it is not * it is dot product

OpenStudy (vishweshshrimali5):

\[\large{\boxed{* \ne \cdot}}\]

OpenStudy (kainui):

(A+B)=2i+0j (A-B)=0i+2j

OpenStudy (anonymous):

yeah .

OpenStudy (vishweshshrimali5):

In vector algebra there are generally two types of products - dot product (scalar product) or cross product (vector product).

OpenStudy (vishweshshrimali5):

If: \[\large{{\bf{U}} = U_x \hat{i} + U_y \hat{j} + U_z \hat{k}}\] \[\large{{\bf{V}} = V_x \hat{i} + V_y \hat{j} + V_z \hat{k}}\] Then: \[\large{{\bf{U}}\cdot {\bf{V}} = U_x V_x + U_y V_y + U_z V_z}\]

OpenStudy (vishweshshrimali5):

This is called dot product ^^^ (The one the problem asks for)

OpenStudy (anonymous):

yess , for the other one we have to make a matrix

OpenStudy (vishweshshrimali5):

Yup that one is vector product

OpenStudy (vishweshshrimali5):

Now that you know how to calculate dot product. Can you find out the answer ?

OpenStudy (anonymous):

yes sure wait :)

OpenStudy (vishweshshrimali5):

Take your time :)

OpenStudy (anonymous):

<2i+0j>.<0i+2j> 0+4-0+0 = 4

OpenStudy (vishweshshrimali5):

How did you get this 4 ?

OpenStudy (vishweshshrimali5):

Why ?

OpenStudy (anonymous):

right**

OpenStudy (vishweshshrimali5):

See: \[\large{\vec{A}\cdot \vec{B} = |\vec{A}||\vec{B}|\cos{\theta}}\] For \(\large{\hat{i},~\hat{j}}\), \(\theta\) is 90 degrees, so dot product is zero.

OpenStudy (vishweshshrimali5):

So i.j = 0 not 1

OpenStudy (vishweshshrimali5):

Thus, answer must be 0 not 4.

OpenStudy (anonymous):

yes 0 is the answer

OpenStudy (vishweshshrimali5):

So do you understand it ?

OpenStudy (anonymous):

Yes really got it (2i).(2j) =0 as i.j=0

OpenStudy (vishweshshrimali5):

Very good :)